A contry's population doubled in 45 yrs. What was the exponential growth rate?
Answer by stanbon(75887)

A contry's population doubled in 45 yrs. What was the exponential growth rate?
-----
2Ao = Ao*2^(t/45)
-----
2 = (2^(1/45))^t
-----
2 = 1.0155^t
----
Growth rate = 1.0155
=========================
Cheers,
Stan H.
=========================